# Evaluation of the upgraded 3-inch Hamamatsu photomultiplier for the KM3NeT Neutrino Telescope

### v2

@@ 1 Introduction
Nominal Voltage for gain $3\times 10^{6}$ 900$\div$1300 V
@@ 3.1 Gain and High Voltage calibration
Figure 6: High voltage tuning results. In the left panel we show the difference between the tuned HV and the values provided by the manufacturer for each PMT. Right: the distribution of the dark count rates (cps) is shown.
The method for determining the HV tuning involves adjusting the high voltage to ensure that the photomultiplier tubes exhibit the same time over threshold (ToT) peak time. This approach is identical to the one used in the study on the R12199-02 [6]. The peak value of the ToT distribution at 26.4 ns for a spe signal at 0.3 spe threshold was used to set the HV value for each PMT, and these values are referred to as the "tuned HV." In Figure 6 (left), the differences between the high voltage (HV) values specified by the vendor and those obtained through tuning in the Dark-Box are shown. This discrepancy arises from different testing conditions: the vendor uses a common resistive voltage divider. In the tests described in this paper the PMTs were equipped with the digital bases used in KM3NeT where the voltage step-up chain is based on a Cockcroft-Walton scheme.
@@ 3.2 Dark count rates
Dark counts can significantly affect the performance of a PMT by generating a background signal that can mask the signals of interest. Therefore, it is important to characterise the dark count rate of PMTs to ensure their proper operation in a detector. The dark count rate is affected by several factors, including the photo cathode area, the photo cathode material, and the natural radioactivity in the structure of the PMT. Bialkali photo cathodes, which are used in KM3NeT PMTs, have a low dark count rate per unit area (100 cps/cm²) wich is common for high-quality, low-noise PMTs . The natural radioactivity in the structure of the PMT can also contribute to the dark count rate, and the most important components are usually 40K and Th contained in the glass envelope. The measured distribution for 1000 PMTs is shown in Figure 6 (right), the mean value obtained is 530 counts per second at the tuned HV value. This result is obtained with a room temperature ideal for human comfort and subject to night day oscillations, however, this the result is well inside the desired specifications for the experiment needs.
@@ 3.3.4 Photoelectron transit time
Transit time and transit time spread are important parameters for time measurements with PMTs. Transit time refers to the time an electron takes to travel from the photo cathode to the first dynode of the PMT, while transit time spread (TTS) refers to the spread in time due to variations in electron trajectories and emission times. In general, smaller TTS values represent for better time resolution in PMTs. Therefore, accurately measuring TTS is important to assess the performance of PMTs and to optimise their use in different experimental settings. To address this the TTS values of PMTs have been measured and compared. The improvement in the time resolution for the R14374 PMT is evident (see right column of Figure 8).
@@ 4 Conclusions
The apparatus described in reference [8] was used to investigate the quantum efficiency of a group of 200 photomultiplier tubes of type Hamamatsu R14374-02. The result indicates that the quantum efficiency curve of the new model is consistent with the one measured in [6]. A radial dishomogeneity through a linear photocatode scan partially documented in the literature [9, 10], was observed. To investigate the radial behaviour over a group of ten R12199-02 and ten R14374-02 PMTs, the circular symmetry of the effect was analysed. The results show better homogeneity in the radial dependence of the quantum efficiency for the new model, as well as a slightly higher QE value in the centre. Also the time properties have been measured. A substantial and evident difference was observed between the two models. The distributions of both afterpulses and delayed pulses favour the new model because the dispersion of the distributions is narrower. The same apparatus previously employed in reference [6] was used to measure the time properties of the new R14374-02 model on a set of 1000 PMTs, and the results showed a noticeable improvement in the time characteristics of this model compared to the previous one. These results suggest that the new model of 3-inch PMT currently in use in the current KM3NeT detectors will improve the overall performance of the entire detector.

### v3

@@ 1 Introduction
Nominal Voltage for gain $3\times 10^{6}$ 900--1300 V
@@ 3.1 Gain and High Voltage calibration
Figure 6: Left: high voltage tuning results. Right: the distribution of the dark count rates (cps) is shown.
The method for determining the HV tuning involves adjusting the high voltage to ensure that the photomultiplier tubes exhibit the same time over threshold (ToT) peak time. This approach is identical to the one used in the study on the R12199-02 [6]. The peak value of the ToT distribution at 26.4 ns for a spe signal at 0.3 spe threshold was used to set the HV value for each PMT, and these values are referred to as the "tuned HV." In Figure 6 (left) the HV values obtained after the tuning process in the Dark-Box are shown.
@@ 3.2 Dark count rates
Dark counts can significantly affect the performance of a PMT by generating a background signal that can mask the signals of interest. Therefore, it is important to characterise the dark count rate of PMTs to ensure their proper operation in a detector. The dark count rate is affected by several factors, including the photocathode area, the photocathode material, and the natural radioactivity in the structure of the PMT. Bialkali photo cathodes, which are used in KM3NeT PMTs, have a low dark count rate per unit area (100 cps/cm²) wich is common for high-quality, low-noise PMTs . The natural radioactivity in the structure of the PMT can also contribute to the dark count rate, and the most important components are usually 40K and Th contained in the glass envelope. The measured distribution for 1000 PMTs is shown in Figure 6 (right), the mean value obtained is 530 counts per second at the tuned HV value. The result was obtained under ambient conditions representative of typical human comfort, with diurnal temperature fluctuations ranging approximately from $20\,^{\circ}\mathrm{C}$ to $25\,^{\circ}\mathrm{C}$. Despite these variations, the outcome remains well within the experimental specifications.
@@ 3.3.4 Photoelectron transit time
Time characteristics of PMTs are measured by detecting and analysing the so called first photon hits, i.e. pulses detected in the window [T0, T0 + 200 ns], where T0 is the calculated arrival time of laser photons on the PMT surface. First hits must have no hits before them in the defined time window. The earliest high peak of the distribution corresponds to the PMT transit time, having already subtracted the travel time of photons in the optical path and the electronic time latency between the laser trigger signal and the laser optical output. These optical and electronics delays are fixed and measured during the DarkBox setting-up [14]. The value of the transit time is determined as the centre of the bin with the maximum counts. The transit time spread is defined as the FWHM of the main peak. It is determined by counting the number of bins for which the content value is greater than half of the maximum bin content. The final distribution of TTS over the PMT set is shown in Figure 8.
@@ 4 Conclusions
The apparatus described in reference [8] was used to investigate the quantum efficiency of a group of 200 photomultiplier tubes of type Hamamatsu R14374-02. The result indicates that the quantum efficiency curve of the new model is consistent with the one measured in [6]. A radial dishomogeneity through a linear photocatode scan partially documented in the literature [9, 10], was observed. To investigate the radial behaviour over a group of ten R12199-02 and ten R14374-02 PMTs, the circular symmetry of the effect was analysed. The results show better homogeneity in the radial dependence of the quantum efficiency for the new model, as well as a slightly higher QE value in the centre. Also the time properties have been measured. A substantial and evident difference was observed between the two models. The distributions of both afterpulses and delayed pulses favour the new model because the dispersion of the distributions is narrower. The same apparatus previously employed in reference [6] was used to measure the time properties of the new R14374-02 model on a set of 1000 PMTs, and the results showed a noticeable improvement in the time characteristics of this model compared to the previous one. The performances described in the manuscript meet the requirements of the experiment and suggests that the new model of 3-inch PMT currently in use in the current KM3NeT detectors will improve the overall performance of the entire detector.